# Fold catastrophe in breaking waves

### v1

@@ 5 Wave breaking following the inflection point
McAllister2023 performed numerical simulations of 22D inviscid and irrotational water waves without surface tension. By tracking the surface point of maximum inflection, where $\eta_{xx}=0$, they found that the minimum local interface slope $m_{b}=\tan(\pi/6)\approx 0.5774$ at that point separates breaking and non-breaking waves. This threshold corresponds to an angle of $30^{\circ}$, matching the limiting steepness derived by Stokes Stokes1847. Such a threshold is not robust, unlike the one following the energetic hotspot. Boettger2025 pointed out that surface tension or wind stresses act as additional forces along the curved air–water interface, affecting the critical slope $m_{b}$ that increases to nearly $60^{\circ}$. Note that the inflection point of a focusing right (left)-going wave is located on the front (back) face of the wave, and its slope is negative (positive) decreasing (increasing) as the wave steepens.
From Eq. (94) in appendix 13), the surface slope $m$ tracked at the inflection point evolves according to the ODE:
@@
The locus $m_{b}^{-}$ is always unstable, whereas $m_{b}^{+}$ is always stable, independent of the value of $m_{2}$, as illustrated in Fig. 7. The two roots are distinct since the discriminant $\Delta=4(m_{2}^{2}+1)$ never vanishes. The line $m_{2}=m$ intersects the two branches at the points $I_{1}$ and $I_{2}$, which correspond to the Stokes limiting slopes $\mp m_{b}$ for right-going (-) and left-going (++) waves, respectively. In the same figure, we show the trajectories (blue lines) of the four tallest successive waves extracted from the numerical simulation of the right-going unidirectional focusing group $C3N5A0.514$ reported in BarthelemyJFM2018. The surface and energy profiles of wave $2$, together with the location of its inflection point, are displayed in Fig. 1.
The four trajectories follow the stable and unstable directions of the dynamical flow; however, a clear breaking inception at $I_{1}$, directly associated with a cusp catastrophe, is not evident. However, breaking occurs for wave ($3$) as the surface slope exceeds the Stokes limit $m_{b}$, in agreement with McAllister2023. At the breaking inception ($I_{1}$), the momentum parameter takes the value $m_{2}=-m_{b}\approx-0.5774$, which is negative and comparable with $m_{2}^{*}=-1$ observed from the energetic hotspot. Thus, the vertical momentum transport is not dominant at the inflection point, and horizontal advection transports momentum toward the crest region.
The physical mechanisms leading to breaking inception at the critical Stokes slope can be identified by examining the horizontal speed $C_{ip}$ of the inflection point, derived in Appendix 12.8:
@@ 6 Black hole analogy and geometrization of wave breaking
We have observed that, as a wave approaches the onset of breaking, its effective dynamics slow down. This is analogous to matter approaching a non-stationary black hole (Vaidya1951). A Vaidya spacetime is a generalized Schwarzschild-like black hole with time-dependent mass, capable of describing collapsing, accreting systems or radiating stars losing mass through light emission (see, e.g., Carroll2004).
The standard gravity analogue is typically based on the characteristic speed of wave perturbation fields of a fluid flow, allowing for a spatially varying speed of sound through the medium (UnruhGravityanalogue2002; GravityAnalogue2011_Liberati). In contrast, the analogy presented here does not rely on any perturbative fluid waves; instead, it emerges directly from the intrinsic, fully nonlinear dynamics of the energetic hotspot of the evolving wave surface.
@@
with explicit forms for specific observables given in Appendix 12. They play the role of matter fields that source the effective geometry. The associated energy-stress tensor can be easily derived from the Einsten equations, which represents the matter and energy required to generate the geometry induced by the evolving wave field.
Null rays, apparent horizon, and event horizon
@@
The apparent horizon is defined as the locus $r=r_{\mathrm{AH}}$ where light rays momentarily stop:
@@
Intuitively, a spaceship located at $r_{\mathrm{AH}}$ sends a light pulse outward; the pulse does not propagate instantaneously outward but hovers at the horizon. Inside the region where $C_{e}<U$, no outward trajectory is possible. This is the gravity analogue of the local breaking onset of the wave. There is no singularity at $r=r_{\mathrm{AH}}$: the Ricci curvature remains finite and null rays cross the apparent horizon smoothly (see, e.g., Carroll2004). The true singularity of the effective geometry occurs where the Ricci scalar diverges.
The event horizon is the global boundary separating light rays that ultimately escape from those that never do. It corresponds to the last null geodesic that reaches distant observers. Unlike the apparent horizon, its location depends on the full future evolution of the wave field.
In gravitational collapse of black holes (the analogue of waves that inevitably break), the event horizon is typically crossed before the apparent horizon. The event horizon is the global boundary beyond which escape is no longer possible, even though nothing catastrophic happens locally at that point. By contrast, the apparent horizon is local where light rays momentarily stop. A key point is that a light pulse crossing the apparent horizon only locally realizes that it can no longer escape, but the causal trap has already begun earlier: the pulse entered the black-hole region when it crossed the event horizon. Thus, the apparent horizon (breaking onset) is where light (wave) “learns too late” that escape is impossible, while the event horizon (breaking inception) marks the true point of no return.
@@ 7 Conclusions
Extending the fast–slow model (26) beyond the breaking region requires a multiple-scale perturbation method to capture higher-order surface-slope effects and to determine the hotspot speed $C_{e}$ in Eq. (85) away from the turning point $O$ (see Fig. 3). In this broader regime, the fold catastrophe may represent a cross-section of a cubic cusp or even a higher-order swallowtail catastrophe (Thom1975; zeeman1976catastrophe; arnold1984catastrophe).
The inception of breaking is identified as associated with the crossing of a well-defined threshold at the characteristic slope angle $\theta^{*}=\arctan{\sqrt{2}-1}=22.5^{\circ}$. This critical angle marks the maximum anisotropy between the Hessians of the velocity and pressure fields that can be sustained by a wave. Physically, this anisotropy reflects an imbalance between the kinetic and potential energy fluxes that govern the local wave dynamics. This trend is robust, as it consistently appears across other simulated breaking wave groups reported in (BarthelemyJFM2018). A similar anisotropy mechanism underlies the emergence of the $30^{\circ}$ slope threshold for breaking inception at the inflection point, beyond which the anisotropy increases. We point out that there is no catastrophe that guides the wave dynamics following the inflection point.
@@ 8 Acknowledgments
The author gratefully acknowledges X. Barthelemy and D. Boettger for generously sharing their numerical data on breaking and non-breaking wave groups. Special thanks are extended to D. Boettger and M. L. Banner for their support and insightful discussions during the preparation of this manuscript. The author thanks Cristel Chandre for revising an early draft of the manuscript and for suggesting the search for asymptotic dynamical models valid near breaking. The author also thanks D. Clamond and D. Dutykh for insightful discussions on steady Stokes waves and group velocities. A special thanks to Arash Yavari, Cristel Chandre and Chongchun Zeng for insightful discussions on geometry and dynamical systems.
@@ 10 Declaration of Competing Interests
11 Data availability
The data that support the findings of this study are available from X. Barthelemy and D. Boettger. AI assistance was limited to language editing and proofreading.
{appen}
12 Time evolution of local observables at a moving surface point
@@
where $R[\eta,\mathbf{U}]$ is a functional of the field variables and their derivatives. The associated velocity vector is $\mathbf{V}_{Q}=(C,\dot{z}_{Q})$ (see, e.g., FedeleEPL2014), where the horizontal component is
@@
with all derivatives evaluated at the moving point $Q$. For example, BarthelemyJFM2018 investigated breaking inception by tracking a wave crest where $R=\eta_{x}=0$, with $C=-\partial_{xt}\eta/\partial_{xx}\eta$ (FedeleEPL2014; FedeleJFMcrestspeed2020). Seliger1968 and McAllister2023 tracked surface points of maximum inflection, where $R=\partial_{xx}\eta=0$ and $C=-\eta_{xxt}/\eta_{xxx}$. In contrast, BoettgerJFM2023; BoettgerPRF2024 followed surface points of maximum kinetic energy.
@@
where all partial derivatives are evaluated at the moving point $Q$. Drawing on Bridges_2009, it is convenient to define the associated one-dimensional (11D) observable
@@ 12.3 Fluid velocities $U$ and $W$
12.4 Kinetic energy density $K_{e}$
@@
12.5 Velocity gradients $U_{x},U_{z}$
For inviscid fluid flows, the Lagrangian derivative of the velocity gradient tensor $\nabla\mathbf{u}=[\partial_{a}u_{b}]$ is (ohkitani1995pressureHessian; Chellard2006pressureHessian)
@@ 12.5 Velocity gradients $U_{x},U_{z}$
12.6 Velocity $C_{e}$ of the energetic hotspot
@@
12.8 Horizontal speed $C_{ip}$ of the inflection point
@@ 13 The governing ODE system
14 The slow-fast ODE system
@@
14.1 Tracking the energetic hotspot
Consider tracking the most energetic point $Q$ on the wave surface, or hotspot, where the kinetic energy density $\tilde{k}_{e}$ is maximum (BoettgerJFM2023). From Eq. (69), the energy gradient $\partial_{x}\tilde{k}_{e}$ is null at the hotspot and defines the momentum parameter in Eq. (15) as
@@ 14.1 Tracking the energetic hotspot
We conjecture that as the wave approaches breaking, its effective dynamics slow down, analogous to nearing a black-hole horizon, suggesting a gravity-analogue interpretation (Faccio2016Emergent). In this view, higher-order pressure-gradient terms follow from imposing next-order slowness in the hierarchical perturbation expansion, without solving the nonlocal pressure equation (13).
15 Equivalent crest speed
@@
The time derivative (FedeleJFM2016)

### v2

@@ 5 Wave breaking following the inflection point
McAllister et al. (2023) performed numerical simulations of 22D inviscid and irrotational water waves without surface tension. By tracking the surface point of maximum inflection, where $\eta_{xx}=0$, they found that the minimum local interface slope $m_{b}=\tan(\pi/6)\approx 0.5774$ at that point separates breaking and non-breaking waves. This threshold corresponds to an angle of $30^{\circ}$, matching the limiting steepness derived by Stokes Stokes (1847). Such a threshold is not robust, unlike the one following the energetic hotspot. Boettger et al. (2025) pointed out that surface tension or wind stresses act as additional forces along the curved air–water interface, affecting the critical slope $m_{b}$ that increases to nearly $60^{\circ}$. Note that the inflection point of a focusing right (left)-going wave is located on the front (back) face of the wave, and its slope is negative (positive) decreasing (increasing) as the wave steepens.
From Eq. (94) in B, the surface slope $m$ tracked at the inflection point evolves according to the ODE:
@@
The locus $m_{b}^{-}$ is always unstable, whereas $m_{b}^{+}$ is always stable, independent of the value of $m_{2}$, as illustrated in Fig. 7. The two roots are distinct since the discriminant $\Delta=4(m_{2}^{2}+1)$ never vanishes. The line $m_{2}=m$ intersects the two branches at the points $I_{1}$ and $I_{2}$, which correspond to the Stokes limiting slopes $\mp m_{b}$ for right-going (-) and left-going (++) waves, respectively. In the same figure, we show the trajectories (blue lines) of the four tallest successive waves extracted from the numerical simulation of the right-going unidirectional focusing group $C3N5A0.514$ reported in Barthelemy et al. (2018). The surface and energy profiles of wave $2$, together with the location of its inflection point, are displayed in Fig. 1.
The four trajectories follow the stable and unstable directions of the dynamical flow; however, a clear breaking inception at $I_{1}$, directly associated with a cusp catastrophe, is not evident. However, breaking occurs for wave ($3$) as the surface slope exceeds the Stokes limit $m_{b}$, in agreement with McAllister et al. (2023). At the breaking inception ($I_{1}$), the momentum parameter takes the value $m_{2}=-m_{b}\approx-0.5774$, which is negative and comparable with $m_{2}^{*}=-1$ observed from the energetic hotspot. Thus, the vertical momentum transport is not dominant at the inflection point, and horizontal advection transports momentum toward the crest region.
The physical mechanisms leading to breaking inception at the critical Stokes slope can be identified by examining the horizontal speed $C_{ip}$ of the inflection point, derived in A.8:
@@ 6 Black hole analogy and geometrization of wave breaking
We have observed that, as a wave approaches the onset of breaking, its effective dynamics slow down. This is analogous to matter approaching a non-stationary black hole (Vaidya, 1951). A Vaidya spacetime is a generalized Schwarzschild-like black hole with time-dependent mass, capable of describing collapsing, accreting systems or radiating stars losing mass through light emission (see, e.g., Misner et al. (1973); Carroll (2004)).
The standard gravity analogue is typically based on the characteristic speed of wave perturbation fields of a fluid flow, allowing for a spatially varying speed of sound through the medium (Schützhold and Unruh, 2002; Barceló et al., 2011). In contrast, the analogy presented here does not rely on any perturbative fluid waves; instead, it emerges directly from the intrinsic, fully nonlinear dynamics of the energetic hotspot of the evolving wave surface.
@@
with explicit forms for specific observables given in A. These fields play the role of effective matter sources that generate the geometry. From the Einstein equations one can formally associate an energy–momentum tensor to this geometry that should satisfy the covariant conservation laws. This condition ensures that energy–momentum is transported consistently along the geometry (Misner et al., 1973; Carroll, 2004).
Null rays, apparent and event horizons
@@
The apparent horizon is defined as the locus $r=r_{\mathrm{AH}}$ where light rays momentarily stop as observed by a distant observer:
@@
Intuitively, a spaceship located at $r_{\mathrm{AH}}$ sends a light pulse outward; the pulse does not propagate instantaneously outward but hovers at the horizon. Inside the region where $C_{e}<U$, no outward trajectory is possible. This is the gravity analogue of the local breaking onset of the wave. There is no singularity at $r=r_{\mathrm{AH}}$: the Ricci curvature remains finite and null rays cross the apparent horizon smoothly. The true singularity of the effective geometry occurs where the Ricci scalar diverges (see, e.g., Misner et al. (1973); Carroll (2004)).
The event horizon is the global boundary separating light rays that ultimately escape from those that never do. It corresponds to the last null geodesic that reaches distant observers. Unlike the apparent horizon, its location depends on the full future evolution of the wave field (Misner et al., 1973; Thornburg, 2007).
In gravitational collapse of black holes (the analogue of waves that inevitably break), the event horizon is typically crossed before the apparent horizon. The event horizon is the global boundary beyond which escape is no longer possible, even though nothing catastrophic happens locally at that point. By contrast, the apparent horizon is local where light rays momentarily stop (Thornburg, 2007). A key point is that a light pulse crossing the apparent horizon only locally realizes that it can no longer escape, but the causal trap has already begun earlier: the pulse entered the black-hole region when it crossed the event horizon. Thus, the apparent horizon (breaking onset) is where light (wave) realizes too late that escape is impossible, while the event horizon (breaking inception) marks the true point of no return.
@@ 7 Conclusions
Extending the fast–slow model (26) beyond the breaking region requires a multiple-scale perturbation method to capture higher-order surface-slope effects and to determine the hotspot speed $C_{e}$ in Eq. (85) away from the turning point $O$ (see Fig. 3). In this broader regime, the fold catastrophe may represent a cross-section of a cubic cusp or even a higher-order swallowtail catastrophe (Thom, 1975; Zeeman, 1976; Arnold, 1992).
The inception of breaking is identified as associated with the crossing of a well-defined threshold at the characteristic slope angle $\theta^{*}=\arctan{\sqrt{2}-1}=22.5^{\circ}$. This critical angle marks the maximum anisotropy between the Hessians of the velocity and pressure fields that can be sustained by a wave. Physically, this anisotropy reflects an imbalance between the kinetic and potential energy fluxes that govern the local wave dynamics. This trend is robust, as it consistently appears across other simulated breaking wave groups reported in (Barthelemy et al., 2018). A similar anisotropy mechanism underlies the emergence of the $30^{\circ}$ slope threshold for breaking inception at the inflection point, beyond which the anisotropy increases. We point out that there is no catastrophe that guides the wave dynamics following the inflection point.
@@ 8 Acknowledgments
The author gratefully acknowledges X. Barthelemy and D. Boettger for generously sharing their numerical data on breaking and non-breaking wave groups. Special thanks are extended to D. Boettger and M. L. Banner for their support and insightful discussions during the preparation of this manuscript. The author thanks Cristel Chandre for revising an early draft of the manuscript and for suggesting the derivation of asymptotic dynamical models valid near breaking. The author also thanks D. Clamond and D. Dutykh for insightful discussions on steady Stokes waves and group velocities. A special thanks to Arash Yavari, Cristel Chandre and Chongchun Zeng for insightful discussions on geometry and dynamical systems.
@@ 10 Declaration of Competing Interests
11 Data Availability Statement
The data that support the findings of this study are available from X. Barthelemy and D. Boettger.
Author Contribution Statement
F. Fedele: Conceptualization, Methodology, Software, Writing - Original Draft.
Appendix A Time evolution of local observables at a moving surface point
@@
where $R[\eta,\mathbf{U}]$ is a functional of the field variables and their derivatives. The associated velocity vector is $\mathbf{V}_{Q}=(C,\dot{z}_{Q})$ (see, e.g., Fedele, 2014a), where the horizontal component is
@@
with all derivatives evaluated at the moving point $Q$. For example, Barthelemy et al. (2018) investigated breaking inception by tracking a wave crest where $R=\eta_{x}=0$, with $C=-\partial_{xt}\eta/\partial_{xx}\eta$ (Fedele, 2014a; Fedele et al., 2020). Seliger (1968) and McAllister et al. (2023) tracked surface points of maximum inflection, where $R=\partial_{xx}\eta=0$ and $C=-\eta_{xxt}/\eta_{xxx}$. In contrast, Boettger et al. (2023, 2024) followed surface points of maximum kinetic energy.
@@
where all partial derivatives are evaluated at the moving point $Q$. Drawing on Bridges (2009), it is convenient to define the associated one-dimensional (11D) observable
@@ A.3 Fluid velocities $U$ and $W$
A.4 Kinetic energy density $K_{e}$
@@
A.5 Velocity gradients $U_{x},U_{z}$
For inviscid fluid flows, the Lagrangian derivative of the velocity gradient tensor $\nabla\mathbf{u}=[\partial_{a}u_{b}]$ is (Ohkitani and Kishiba, 1995; Chevillard and Meneveau, 2006)
@@ A.5 Velocity gradients $U_{x},U_{z}$
A.6 Velocity $C_{e}$ of the energetic hotspot
@@
A.8 Horizontal speed $C_{ip}$ of the inflection point
@@ Appendix B The governing ODE system
Appendix C The slow-fast ODE system
@@
C.1 Tracking the energetic hotspot
Consider tracking the most energetic point $Q$ on the wave surface, or hotspot, where the kinetic energy density $\tilde{k}_{e}$ is maximum (Boettger et al., 2023). From Eq. (69), the energy gradient $\partial_{x}\tilde{k}_{e}$ is null at the hotspot and defines the momentum parameter in Eq. (15) as
@@ C.1 Tracking the energetic hotspot
We conjecture that as the wave approaches breaking, its effective dynamics slow down, analogous to nearing a black-hole horizon, suggesting a gravity-analogue interpretation (Marino et al., 2016). In this view, higher-order pressure-gradient terms follow from imposing next-order slowness in the hierarchical perturbation expansion, without solving the nonlocal pressure equation (13).
Appendix D Equivalent crest speed
@@
The time derivative (Fedele et al., 2016)